The sum of Al and Bill is 25, the sum of the ages Al and Carl is 29, the sum of the ages of Billa and Car is 32. Who is the oldest of the three and how old is he?
Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

A+B=25
A+C=29
B+C=32
C=32-B
therefor A+C=A+32-B=29 or A-B=-3
A+B=25
add them
2A=22
A=11
B=14
C=18

Answer by greenestamps(13200)



If the problem were only to determine who is oldest, then logical reasoning immediately tells us it is Carl.

To find Carl's age, I would use a different strategy than the other tutor.

Given:
A+B = 25
A+C = 29
B+C = 32

Add the three equations:
2(A+B+C) = 86
A+B+C = 43

Carl's age is the sum of all three ages, minus the sum of Al's and Bill's ages.
C = (A+B+C)-(A+B) = 43-25 = 18.

The ages of the others can be found, if needed, in a similar way.